Sam has scored 14 ,20, 22, and 16 points in 4 basket ball games how many points must he score in the 5th game .if the average is to be 20 ?
A 32 B 28 C 24 D 22 E 20

Answer by AAfter Search(61)

Let the points scored in the fifth game be x for the average to be 20.
Given, (14 + 20 + 22 + 16 + x) / 5 = 20
=> (72 + x) / 5 = 20
=> 72 + x = 100
=> x = 100 - 72 = 28
Hence, Sam must score 28 points in the fifth game to have an average of 20.